TPC QA for run 2023

July 3  run 184006
Fixes after daq5K
Tonko: Unmasked a series of iTPC RDOs: 
iS9-1, iS9-3, iS9-4 
iS24-1 iS24-4 

4 RDOs are still masked: 
iS06-1 doesn't work at all with daq5k (stays masked until the end of the run) 
iS11-3 likely bad power (Anderson) connector -- stays masked iS17-3 low voltage on input -- waiting for Wed access 
iS18-4 still needs work on my end 
iS19-2 still needs work (recently masked)

may 4 2023
The startup for cosmics has been dismal.
  1. Trigger issues - resulted in many days down time. TCU replacement
  2. 6 Hz limit for some TPC readout. Not clear how this was resolved. Masking out many iTPC sectors helped.
  3. Waterleak
As of today we may be able to have cosmics back late Friday

Status of iTPC FEEs as per Tonko 5/4/2023

  • iTPC bad FEEs on 04-May-2023 (superceeds all below)

    • Sector RDO Port …

    • 1 2 3 -1 0 0 0.00 0.00

    • 8 4 8 -1 0 0 0.00 0.00

    • 11 4 8 -1 0 0 0.00 0.00

    • 9 4 4 -1 0 0 0.00 0.00

    • 2 1 8 -1 0 0 0.00 0.00

    • 7 4 6 -1 0 0 0.00 0.00 # 27Apr23, POK

    • 11 2 14 -1 0 0 0.00 0.00 # 30Apr23: syncs but doesn't configure

    • 20 3 12 -1 0 0 0.00 0.00     # 04May23: corruption in data...

May 5, 2023
status update from Tonko after 2. water leak checks

We are having problems with 1 iTPC RDO (iS11-3)
starting before the poletip insertion (but after removal).
It's power is flaky but I need to re-check with Christian
if we can fix it externally.

Also, one iTPC FEE in Sector 7 (iS7-2-5) is not responding.
Looks like cable problems so potentially related to
poletip installation. It was fine yesterday. But I'm not sure if 1 FEE warrants
another poletip removal/installation.

May 6, 2023

There was a succesfull laser run 126046
The peculir think is there are no traces of the tracks in the 2 outer 5,6 RDO in TPX, even though the ADC
plots from online shows this clearly. It was mentioned at 10 o'clock that the gains may not have been applied properly.
that the reason here.

This is the cluster plots (per event)

The issue from above was fixed after several days of work by tonko and Jfee. It was a data bank naming issues dues to multiple PC per TPX sector.

First observation of water damage
yesterday after the magnet came back on and on the
first run with TPC (6:50PM) a large number of TPC
RDOs failed. ALL those RDOs are in Sectors 4, 5 & 6
and they are different electronics, iTPC & TPX.

This morning (now) some of those RDOs seem to have
recovered. The current state is:

1) Sector 5: ALL 4 inner (iTPC) RDOs are dead for various
reasons. ALSO outer sector (TPX) RDO 3 is dead.
In other words a progression of different electronics from inner
to outer radius is shot.

2) Sector 4: Inner ITPC RDO1 is dead. Note that this is
the innermost piece of electronics.

3) Sector 6 is all fine, inner & outer, recovered.

It looks pretty obvious that something localized to that area happened.
One can suspect a water leak. Especially since some RDOs have
recovered since yesterday perhaps due to drying out.
Or one can suspect something moved related to the B-field going off
and on. I have no way of telling.

I shut the power down on  sectors 4, 5 & 6, inner & outer, just in
case of a water leak.
I will suggest to the Shifcrew not to run iTPC & TPX until we discuss
this at 10AM to avoid confusion with masking/unmasking various
sectors/RDOs etc.etc.

5/14/2023  Likely water issue lost many ROD's in sector 4,5 and 6
Here is Tonko's summary of proposed priority for repairs on 5/16/23

for those who want details about the failed RDOs:

Priority 1) Critically damaged RDOs: 4 inner, 1 outer
Priority 2) RDOs with significant count of bad FEE ports: 3 inner
Priority 3) RDOs with smaller number of bad FEE ports: 2 outer

All are in Sectors 4,5 or 6 as before with Sector 5 being the worst
(damage to 6 RDOs)
followed by Sector 6 (damage to 3 RDOs) and then Sector 4 (just 1 RDO).

This is my proposed order of swapping priority by RDO:

Priority 1: iS5-2, iS5-4, iS6-2, iS6-4, oS6-3
Priority 2: iS5-1, iS5-3, iS4-1
Priority 3: oS5-3, oS5-4

We will do a quick check of the RDOs with the failed FEE ports
by swapping fee cables to make sure it is on the RDO side and not the FEE.
If it's the FEE we will try to swap (some of) those FEEs but some of
the FEEs might
not be easily reachable without further disassembly so we'll make
the decision on the spot.